OpenStudy (anonymous):

Is it possible to have zero acceleration but non-zero velocity at some instant in a motion?

OpenStudy (agreene):

Yes. If net force is zero.

OpenStudy (anonymous):

what are net force? What do you mean?

OpenStudy (agreene):

Net Force is the sum of forces acting on a body. Since: \[F_{NET}=m*a\] and \[F_{NET}\alpha A\] we can say that if Net Force is 0, Acceleration is zero. And, because we know: R is the displacement vector, and: R'(t) = V(t) R''(t) = A(t) we can suppose a case where: A(t) = 0 and find our velocity from: \[\int A(t) dt=\int0dt=C_{1}\] where C is our integration constant. meaning that our velocity is some constant value. From this we can also find the displacement from: \[\int V(t) dt=\int C_{1}dt= C_{1}t+C_{2}\] making this a linear equation for R(t) which makes sense... constant velocity = linear displacement.

OpenStudy (anonymous):

I'm so sorry but I have no idea what you just put above D:

OpenStudy (agreene):

Basically, the answer is yes. It happens when the sum of all the forces acting on an object is 0. Because the relation between Force and acceleration, and acceleration and velocity. I used calculus to show this, because it is quicker and easier than any other method.

OpenStudy (jamesj):

The short answer is yes because force leads to acceleration of an object. However, if the object is already moving--i.e., has some velocity--then with no force acting on it, it will continue to keep moving. Perhaps the easiest analogy is on a billiard table, with negligible friction. You hit a ball; it accelerates; then it starts to move with constant velocity even though there no longer any force applied to it.

OpenStudy (stormfire1):

You should consider reviewing Newton's First Law (law of inertia): http://science.howstuffworks.com/science-vs-myth/everyday-myths/newton-law-of-motion1.htm

OpenStudy (anonymous):

yes, it's possible.. if the velocity is constant , the acceleration is zero.